OpenStudy (anonymous):

WILL GIVE MEDAL Use the table below to answer the question. x y 0 2 1 -1 2 -6 Find the average rate change for the given function form x=0 to x=2

OpenStudy (anonymous):

@iGreen

OpenStudy (studygurl14):

Find the slope using points (0,2) and (2,-6)

OpenStudy (studygurl14):

Do you know the slope formula?

OpenStudy (anonymous):

-4

OpenStudy (igreen):

Our two points are (0, 2) and (2, -6) We can plug these into the slope equation: \(m = \dfrac{y_2-y_1}{x_2-x_1}\) \(m = \dfrac{y_2-y_1}{x_2-x_1}\) \(m = \dfrac{-6-2}{2-0}\) Subtract: \(m = \dfrac{-8}{2}\) Can you divide -8 / 2? @cerickson2

OpenStudy (igreen):

You got it! @cerickson2

OpenStudy (anonymous):

Awesome! can you help me with one more?

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Response'. @cerickson2

OpenStudy (igreen):

Sure, just close this question and open a new one.

OpenStudy (anonymous):

oogaly boogaly